# Uncovering Grounding IDs: How External Cues Shape Multi-Modal Binding

### v2

@@ 4.1 Activation Swapping Experiment
Next, we investigate in which layers the model begins to predict the intervened grounded object ($\textbf{o}^{s}_{\sim s}$) rather than the adjacent object ($\textbf{o}^{\sim s}_{s}$). We apply the logit lens technique (nostalgebraist, 2020) to the single-word response token (using a monochrome dataset for simplicity). By computing the logit difference between the $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ and plotting averaged values across layers (Fig. 5(a)), we find that the model shifts toward predicting the intervened grounded object in the later layers (20–27). We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\textbf{o}^{s}_{\sim s}$ versus $\textbf{o}^{\sim s}_{s}$ across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers grounded objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in layer 16, which attends more to grounded objects.
@@ 4.2 The Characteristics of Grounding IDs
Feng & Steinhardt (2023) shows that in LLMs, Binding IDs linking entities and their attributes are largely context independent. In contrast, our causal mediation experiment demonstrates that Grounding IDs are directly predictable from their corresponding symbols. To probe this relationship, we analyze the relational structure of symbols and their induced Grounding IDs. Concretely, for each pair of symbols (e.g., &, #), we compute the difference between their symbol patch activations. For the same pairs, we also compute the difference between the corresponding object patch activations. Averaging these differential vectors across the dataset cancels out confounding factors such as shape and position, yielding two structured spaces: one defined by symbol differences, and one by Grounding ID differences. Fig. 6 reports cosine similarities between these two spaces, revealing strong correspondence between the symbol space and the Grounding ID space.
Figure 6: Cosine similarity between averaged differential vectors of symbol patches and their corresponding Grounding IDs.
@@
5 Implications of Grounding IDs: Hallucination Reduction
@@
Figure 7: Cross-attention behavior on the synthetic dataset with 20 objects, averaged over 100 samples. Attention is computed using a max operation with a window size of 5 and stride of 1 across generated tokens.
Table 1: Evaluation of the scene description task on synthetic datasets, each containing 100 samples with 10, 15, or 20 unique objects per image.
# Objects Method Precision Recall F1 Accuracy
10 Baseline 0.56 0.60 0.58 0.41
Structured (text-only) 0.57 0.61 0.59 0.42
Structured (image-only) 0.60 0.56 0.58 0.40
@@
15 Baseline 0.30 0.49 0.37 0.22
Structured (text-only) 0.37 0.53 0.50 0.27
Structured (image-only) 0.40 0.47 0.43 0.27
Structured 0.67 0.53 0.59 0.42
20 Baseline 0.18 0.47 0.26 0.15
Structured (text-only) 0.32 0.45 0.37 0.23
Structured (image-only) 0.40 0.46 0.43 0.26
Structured 0.67 0.50 0.57 0.40
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (Favero et al., 2024; Huang et al., 2024). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, measuring accuracy, precision, recall, and F1 for object descriptions, as reported in Table 1. Structured inputs consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens with increasing object counts. Notably, multimodal cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (Lin et al., 2014), with performance assessed by the hallucination-specific metric CHAIR (Rohrbach et al., 2018). We conduct experiments on two recent LVLMs, Qwen2.5-VL (Bai et al., 2025), and LLaVA-1.5 (Liu et al., 2023a). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding. Results (Fig. 8 and Table 2) show substantial reductions in both CHAIRs and CHAIRi compared to baseline. Importantly, unlike most existing techniques, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (Hurst et al., 2024) and Gemini-2.5-Pro (Comanici et al., 2025), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and SPARC (Jung et al., 2025), while requiring no additional inference modules and maintaining near-zero computational overhead. We also report results on another commonly used benchmark, POPE (Li et al., 2023), in the appendix.
@@
Table 2: Evaluation on MS-COCO using CHAIR metrics across open- and closed-source models. Results are reported for sentence-level (CHAIRS) and instance-level (CHAIRI) hallucination rates (lower is better). The average inference time measured on a machine with an NVIDIA RTX 4090 GPU.
@@
LLaVA-1.5 Baseline 59.00 13.40 3.41
Opera 49.50 11.90 20.91
VCD 49.50 11.78 7.81
SPARC 61.00 11.93 4.50
Structured 39.00 11.43 3.94
Qwen2.5-VL Baseline 40.00 8.45 3.31
Opera 28.00 7.40 23.50
VCD 35.50 8.05 9.73
SPARC 38.50 8.45 5.50
Structured 28.50 4.57 6.04
GPT-4o Baseline 29.00 6.41 -
Structured 28.50 4.97 -
Gemini-2.5-Pro Baseline 49.50 8.69 -
Structured 40.00 7.36 -
@@ 6 Related Work
Interpretability research investigates where and how information is represented and routed in neural networks. For LLMs, attention analyses study which tokens heads attend to and how this shapes predictions (Chefer et al., 2021; Clark et al., 2019), while circuit-oriented methods identify causal substructures and enable editing through activation or circuit discovery and factual localization (Conmy et al., 2023; Meng et al., 2022; Geva et al., 2023). Probing tools such as the logit lens decode intermediate states to reveal how token predictions evolve across layers (nostalgebraist, 2020). For VLMs, recent work explores where visual information is stored and how it is transferred into the language pathway (Basu et al., 2024), and adapts logit-lens–style spatial probes to study grounding and localization inside multimodal models (Jiang et al., 2024; Neo et al., 2024). Overall, these methods provide the toolset we build on: causal interventions and layerwise probes to expose mechanisms underlying cross-modal binding and grounding.
@@
Recent advances suggest that models internally rely on abstract latent variables to maintain entity–attribute associations. In LLMs, Binding IDs are content-independent identifiers that link entities and attributes through a shared latent code; causal interventions that swap these vectors systematically change the inferred associations (Feng & Steinhardt, 2023). Follow-up work localizes a low-rank subspace that causally governs which entity pairs with which attribute (Dai et al., 2024). In multimodal models, analogous identifiers appear in VLMs: distinct binding codes are attached to an object’s image tokens and its textual mentions, yielding cross-modal referential consistency (Saravanan et al., 2025). Complementary evidence points to symbolic indexing in VLMs, where attention heads compute content-independent spatial indices and later retrieve attributes by these indices, thereby implementing object-centric binding across modalities (Assouel et al., 2025). These findings highlight the role of hidden symbolic variables in reasoning. While prior analyses primarily examine existing identifiers, we show that simple input-augmented artifacts can elicit partition-specific identifiers that propagate across modalities and enhance grounding through tighter cross-modal attention and alignment.
@@
Efforts to improve reasoning in LVLMs have followed two main directions. Prompt-focused methods such as chain-of-thought prompting, majority voting, and test-time compute scaling aim to enhance reasoning without altering inputs. However, evaluations on multimodal reasoning benchmarks show that these approaches remain ineffective for tasks requiring visual or spatial reasoning, even when scaled substantially (Hao et al., 2025). In contrast, recent studies demonstrate that external scaffolding—adding annotations, grids, or partitions to the input image—consistently improves model performance across counting, spatial relations, and description tasks (Rudman et al., 2025; Izadi et al., 2025). While such findings establish the effectiveness of scaffolding, the mechanisms underlying these improvements remain poorly understood.
@@
Hallucination remains a central challenge for LVLMs, mainly arising from the decline of cross-modal attention during long text generation and over-reliance on language priors. Several inference-time methods have been proposed, including VCD (Leng et al., 2024), which down-weights tokens favored under distorted inputs; OPERA (Huang et al., 2024), which applies an over-trust penalty and retrospection allocation to rebalance attention; and SPARC (Jung et al., 2025), which progressively recalibrates visual attention to maintain relevance in long captions. Our work complements these approaches by identifying how visual scaffolding induces Grounding IDs that naturally sustain cross-modal binding, thereby reducing hallucination without additional inference tools.
@@ 7 Conclusion
In this work, we introduce a conceptual framework to explain how external cues induce abstract Grounding IDs across related partitions. Through attention and embedding alignment analysis, we showed that these identifiers propagate across modalities, supporting partition-specific grounding and reducing the modality gap. Activation patching interventions confirmed that the associations are mediated by abstract identifiers rather than local features. Empirical evaluations further demonstrated that Grounding IDs enhance cross-modal attention, yield more faithful image descriptions, and reduce hallucinations. Since the approach relies only on simple, content-independent structures, it provides a model-agnostic strategy applicable to a wide range of tasks and models, including closed-source LVLMs.
Beyond these empirical contributions, our findings shed light on interpretability in multimodal models by identifying a symbolic mechanism underlying visual reasoning. This opens several directions for future research, particularly in circuit discovery. One is to explore mechanisms supporting system-2 reasoning tasks, such as counting and spatial reasoning. Another stems from a key observation of this study: external cues in VLM inputs can reinforce the model’s inherent grounding capability. Building on this, future work could integrate such cues during RL finetuning to further strengthen the model’s internal ability to perform sequential scanning based on provided cues. Overall, this work bridges mechanistic understanding and practical utility, showing how induced identifiers can both clarify internal dynamics and enhance robustness in LVLMs.
@@ A.5 Implementation Details and Qualitative Example
A.6 Evaluation Metrics for Object Hallucination
@@ A.6.1 CHAIR: Caption Hallucination Assessment with Image Relevance
We compute CHAIR scores on 200 randomly selected images from the MSCOCO validation set, using model-generated captions and official COCO object annotations as references.

### v3

@@ 4.2 Layerwise Analysis of Grounding ID Emergence
We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\textbf{o}^{s}_{\sim s}$ versus $\textbf{o}^{\sim s}_{s}$ across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers grounded objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. Let $\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o})$ denote the attention weight from the response token $\mathbf{r}$ to the patch of object $\mathbf{o}$ at layer $\ell$, head $h$. For each head, we compute
$S^{(\ell,h)}\;=\;\frac{\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}{\operatorname{Std}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}\cdot\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\text{image})\right],$ (2)
where expectations are taken over samples and valid symbol–object pairs. Heads with high $S^{(\ell,h)}$ consistently prefer grounded objects over merely co-located ones and are interpreted as key carriers of Grounding IDs. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in middle layers (layer 16), which attend more to grounded objects.
@@ 4.3 The Characteristics of Grounding IDs
Figure 6: Characteristics of Grounding IDs. (a) Cosine similarity between averaged differential vectors of symbol patches (vertical) and their corresponding Grounding IDs (horizontal). (b) Disjoint-symbol control for the activation-swap test. The source context contains the symbols {&,\$,#,@}, while the target context contains a nonoverlapping set $\times$, +}.
binding_id shows that, in LLMs, Binding IDs linking entities and their attributes are largely context independent. In contrast, our causal mediation experiment demonstrates that Grounding IDs are directly predictable from their corresponding symbols. To probe this relationship, we analyze the relational structure of symbols and their induced Grounding IDs. Concretely, for each pair of symbols (e.g., &, #), we compute the difference between their symbol patch activations. For the same pairs, we also compute the difference between the corresponding object patch activations. Averaging these differential vectors across the dataset cancels out confounding factors such as shape and position, yielding two structured spaces: one defined by symbol differences and one by Grounding ID differences. Fig. 6(a) reports cosine similarities between these two spaces, revealing strong correspondence between the symbol space and the Grounding ID space.
@@
5 Behavioral Implications of Grounding IDs
Our observations in Section 3 exhibit a reduced modality gap and increased attention between related partitions in samples with external cues. The causal analysis in Section 4 further reveals the presence of latent identifiers that bind to the embeddings of the corresponding partitions. Together, these results describe a mechanism by which external cues improve visual reasoning: they slightly modify the model’s internal dynamics and equip visual embeddings with auxiliary identifiers that strengthen the model’s grounding ability. To examine how this property affects downstream behavior, we first evaluate the impact of Grounding IDs on hallucination mitigation during long caption generation, since this task directly depends on grounding. We then assess the effect of aligned multimodal cues on broader visual reasoning tasks.
5.1 Hallucination Mitigation
@@
Figure 7: Averaged cross-attention behavior on the synthetic dataset with 20 objects. Attention is computed using a max operation with a window size of 5 and stride of 1 across generated tokens.
Table 1: Evaluation of the scene description task on synthetic datasets, each containing 500 samples with 10, 15, or 20 unique objects per image.
# Obj. Method Precision Recall F1 Accuracy
10 Baseline 0.56 0.56 0.58 0.42
Structured (text-only) 0.59 0.68 0.63 0.46
Structured (img-only) 0.53 0.59 0.56 0.38
@@
15 Baseline 0.30 0.49 0.37 0.24
Structured (text-only) 0.33 0.61 0.44 0.27
Structured (img-only) 0.43 0.51 0.46 0.30
Structured 0.67 0.53 0.59 0.46
20 Baseline 0.14 0.45 0.21 0.12
Structured (text-only) 0.29 0.57 0.39 0.24
Structured (img-only) 0.39 0.42 0.40 0.24
Structured 0.65 0.59 0.62 0.40
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (m3id; opera). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, measuring accuracy, precision, recall, and F1 for object descriptions, as reported in Table 1. Structured inputs consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens as the number of objects increases. Notably, multimodal cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (ms_coco), with performance assessed by the hallucination-specific metric CHAIR (chair). We conduct experiments on two recent LVLMs, Qwen2.5-VL (qwen2.5_vl), and LLaVA-1.5 (llava1.5). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding.
Results (Fig. 8 and Table 2) show substantial reductions in both CHAIRs and CHAIRi compared to baseline. Importantly, unlike most existing techniques for hallucination mitigation, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (gpt4o) and Gemini-2.5-Pro (gemini2.5), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (vcd), OPERA (opera), and SPARC (sparc), while requiring no additional inference modules and maintaining near-zero computational overhead. We also report results on another commonly used benchmark, POPE (pope), in Appendix A.6
@@
Table 2: Evaluation on 500 MS-COCO samples using CHAIR metrics across open- and closed-source models. Results are reported for sentence-level (CHAIRs) and instance-level (CHAIRi) hallucination rates.
@@
LLaVA-1.5 Baseline 51.60 13.20 3.41
Opera 48.00 13.52 20.91
VCD 54.40 14.28 7.81
SPARC 55.20 12.78 4.50
Structured 41.00 12.04 3.94
Qwen2.5-VL Baseline 32.40 7.97 3.31
Opera 29.60 10.76 23.50
VCD 33.80 8.91 9.73
SPARC 33.60 8.21 5.50
Structured 27.20 5.36 6.04
GPT-4o Baseline 29.20 6.40 -
Structured 23.20 5.81 -
Gemini2.5-Pro Baseline 44.20 8.64 -
Structured 37.40 7.28 -
5.2 Visual Reasoning Performance
We evaluate Grounding IDs on two visual reasoning benchmarks related to spatial grounding: counting and visual search. We conduct experiments on the datasets introduced in viser. The counting dataset consists of synthetic 2D scenes with 10 to 20 circles, increasing in steps of 2, with 100 images for each count level (600 total). The visual search dataset contains conjunctive search scenes with 20, 30, 40, and 50 objects. For each difficulty level, 100 images are created: 50 with the target present and 50 absent. Accuracy is computed separately for visible and invisible cases and combined using the harmonic mean.
We evaluate three models: Qwen-3B, Qwen-7B, and GPT-4o. Grounding IDs increase accuracy on both tasks and outperform the VISER baseline and the unstructured input. The gains depend on two conditions: each partition must have a distinct identifier, and the same identifiers must appear in the image and the prompt. The results are shown in Table 3.
Table 3: Performance on counting and visual search benchmarks.
Counting Accuracy Visual Search Accuracy
Model Baseline VISER Grounding IDs Baseline VISER Grounding IDs
Qwen2.5-VL (3B) 30.00 37.83 43.00 0.00 37.83 45.96
Qwen2.5-VL (7B) 29.67 43.33 53.00 30.00 40.00 52.25
GPT-4o 10.50 26.50 32.33 49.41 73.40 80.62
@@ 6 Related Work
Interpretability research investigates where and how information is represented and routed in neural networks. For LLMs, attention analyses study which tokens heads attend to and how this shapes predictions (chefer21_generic_attention; clark19_bert_attention), while circuit-oriented methods identify causal substructures and enable editing through activation or circuit discovery and factual localization (conmy23_automated_circuit_discovery; meng22_rome_factual_edit; geva23_dissect_recall). Probing tools such as the logit lens decode intermediate states to reveal how token predictions evolve across layers (nostalgebraist20_logit_lens). For VLMs, recent work explores where visual information is stored and how it is transferred into the language pathway (basu24_storage_transfer_vlm), and adapts logit-lens–style spatial probes to study grounding and localization inside multimodal models (jiang24_vl_rep_edit; llava_interp). Overall, the interpretability methods provide the toolset we build on: causal interventions and layerwise probes to expose mechanisms underlying cross-modal binding and grounding.
@@
Recent advances suggest that models internally rely on abstract latent variables to maintain entity–attribute associations. In LLMs, Binding IDs are content-independent identifiers that link entities and attributes through a shared latent code; causal interventions that swap these vectors systematically change the inferred associations (binding_id). Follow-up work localizes a low-rank subspace that causally governs which entity pairs with which attribute (dai24_binding_order). In multimodal models, analogous identifiers appear in VLMs: distinct binding codes are attached to an object’s image tokens and its textual mentions, yielding cross-modal referential consistency (vision_binding_id). Complementary evidence points to symbolic indexing in VLMs, where attention heads compute content-independent spatial indices and later retrieve attributes by these indices, thereby implementing object-centric binding across modalities (visual_symbolic_mechanism). These findings highlight the role of hidden symbolic variables in reasoning. While prior analyses primarily examine existing identifiers, we show that simple input-augmented artifacts can elicit partition-specific identifiers that are linked across modalities and enhance grounding through tighter cross-modal attention and alignment.
@@
Efforts to improve reasoning in LVLMs have followed two main directions. Prompt-focused methods such as chain-of-thought prompting, majority voting, and test-time compute scaling aim to enhance reasoning without altering inputs. However, evaluations on multimodal reasoning benchmarks show that these approaches remain ineffective for tasks requiring visual or spatial reasoning, even when scaled substantially (emma). In contrast, recent studies demonstrate that external scaffolding—adding annotations, grids, or partitions to the input image—consistently improves model performance across counting, spatial relations, and description tasks (forgotten_polygons; viser). While such findings establish the effectiveness of scaffolding, the mechanisms underlying these improvements remain poorly understood.
@@
Hallucination remains a central challenge for LVLMs, mainly arising from the decline of cross-modal attention during long text generation and over-reliance on language priors. Several inference-time methods have been proposed, including VCD (vcd), which down-weights tokens favored under distorted inputs; OPERA (opera), which applies an over-trust penalty and retrospection allocation to rebalance attention; and SPARC (sparc), which progressively recalibrates visual attention to maintain relevance in long captions. Our work complements these approaches by identifying how visual scaffolding induces Grounding IDs that naturally sustain cross-modal binding, thereby reducing hallucination without additional inference tools.
@@ 7 Conclusion
In this work, we introduce a conceptual framework to explain how multimodal aligned external cues induce abstract Grounding IDs across related partitions. Through attention and embedding alignment analysis, we showed that these identifiers propagate across modalities, supporting partition-specific grounding and reducing the modality gap. Activation patching interventions confirmed that the associations are mediated by abstract identifiers rather than local features. Empirical evaluations further demonstrated that Grounding IDs enhance cross-modal attention, yield more faithful image descriptions, and reduce hallucinations. Since the approach relies only on simple, content-independent structures, it provides a model-agnostic strategy applicable to a wide range of tasks and models, including closed-source LVLMs.
Beyond empirical contributions, our methodology provides insights into mechanistic interpretability in multimodal models through both observational and causal tools. In addition, this study opens several directions for future research. One direction involves discovering circuits that support system-2 reasoning tasks, such as counting and spatial reasoning. Another arises from a key observation in this study: external cues in VLM inputs can reinforce the model’s inherent grounding capability. Building on this, future work could explore integrating such cues during RL finetuning to further strengthen the model’s internal ability to perform sequential scanning based on provided cues.
@@ A.4 Evaluation on Synthetic Variants
  • Symbols (no line): Each row is tagged with a unique symbol (&, #, @, \$), but no horizontal separators are drawn. This already improves both precision and accuracy compared to the baseline.
  • Object Bounding (input-aware): Each object is enclosed within an explicit bounding box. While this provides a localized spatial cue, it is less effective than row-level structuring, resulting in only moderate improvements over the baseline.
  • Symbols on Objects (input-aware): Symbols are assigned at the object level instead of the row level. While this introduces additional visual markers, the lack of global row structure leads to reduced performance compared to row-based variants.
@@
Symbols (no line) 0.48 0.50 0.49 0.33
Object Bounding Box 0.43 0.48 0.44 0.32
Symbols on Objects 0.30 0.38 0.29 0.19
@@ A.5 Implementation Details and Qualitative Example
A.6 Evaluation Benchmarks for Object Hallucination
@@ A.6.1 CHAIR: Caption Hallucination Assessment with Image Relevance
We compute CHAIR scores on 500 randomly selected images from the MSCOCO validation set, using model-generated captions and official COCO object annotations as references.